# Towards Causal Market Simulators

### v1

@@ Abstract.
copyright: nonejournalyear: 2025conference: Workshop on Rethinking Financial Time Series; November 15–18, 2025; Singaporebooktitle: 6th ACM International Conference on AI in Finance (ICAIF ’25): Workshop on Rethinking Financial Time Series, November 15–18, 2025, Singaporeisbn: 978-8-4007-2220-2/2025/11ccs: Computing methodologies Artificial intelligenceccs: Applied computing Enterprise financial systemsccs: Computing methodologies Machine learning
@@ 2. Methodology
We propose a causal market simulator building on top of previous research in neural causal models (NCMs) (xia2023neural) and variational autoencoder (VAE) (kingma2013auto) based causal representation learning (scholkopf2021toward) in time series. Our neural causal model for time series (TNCM-VAE) consists of three main components: an encoder to infer latent representations, a causal mapping module to handle dependencies, and a decoder to generate counterfactual sequences. As visualized in Figure 1, $\{X,Y\}\subseteq V$ represents the sets of all windows of time.

### v2

@@ Abstract.
copyright: ccjournalyear: 2025doi: 10.48550/arXiv.2511.04469conference: Workshop on Rethinking Financial Time Series; November 15–18, 2025; Singaporebooktitle: 6th ACM International Conference on AI in Finance (ICAIF ’25): Workshop on Rethinking Financial Time Series, November 15–18, 2025, Singaporeisbn: 978-8-4007-2220-2/2025/11ccs: Computing methodologies Artificial intelligenceccs: Applied computing Enterprise financial systemsccs: Computing methodologies Machine learning
@@ 2. Methodology
We propose a causal market simulator building on top of previous research in neural causal models (NCMs) (xia2023neural) and variational autoencoder (VAE) (kingma2013auto) based causal representation learning (scholkopf2021toward) in time series. Our neural causal model for time series (TNCM-VAE) [^1] https://github.com/thummd/tncm consists of three main components: an encoder to infer latent representations, a causal mapping module to handle dependencies, and a decoder to generate counterfactual sequences. As visualized in Figure 1, $\{X,Y\}\subseteq V$ represents the sets of all windows of time.